OpenStudy (anonymous):

A circle has a circumference of 14 ft. Its area is halved. Describe the effect on the circumference of the circle. The circumference is divided by √2. The circumference is divided by 2. The circumference doubles. The circumference is divided by √3.

OpenStudy (anonymous):

PLEASE I NEED HELP!!!!!!

OpenStudy (anonymous):

Find the surface area of a sphere with a diameter of 76 cm. 3850 π cm2 1925 π cm2 5776 π cm2 7701 π cm2

OpenStudy (tkhunny):

Area Proportion: a^2:b^2 Linea Measure Proportion: a:b Halving the area gives these proportions: a^2:(1/2)b^2 Moving to Linear Measure \(a:\dfrac{b}{\sqrt{2}}\)

OpenStudy (anonymous):

I posted the new question that I need help with :) @tkhunny

OpenStudy (anonymous):

@Squirrels

OpenStudy (tkhunny):

You're not getting the idea posting new question on new threads. Diameter = 2* Radius Volume = \(\dfrac{4}{3}\pi r^{3}\) Surface Area = \(4\pi r^{2}\)

OpenStudy (anonymous):

so how do i solve for the answers?

OpenStudy (anonymous):

Find the surface area of a sphere with a diameter of 76 cm. 3850 π cm2 1925 π cm2 5776 π cm2 7701 π cm2

OpenStudy (anonymous):

those are the options

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

I hate to say this but i am in a awful hurry, my class ends in 2 hours

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

@superhelp101

OpenStudy (anonymous):

@Squirrels

OpenStudy (squirrels):

Just use the surface area formula \[4\pi r^2\]

OpenStudy (squirrels):

Thats the surface area formula for the "sphere".

OpenStudy (anonymous):

how do i do that? im seriously bad at math! :(

OpenStudy (squirrels):

Thats fine, ill teach you how to do it but I'm not giving you the answer just like that :p

OpenStudy (anonymous):

I do really appreciate the help, im just in a hurry cause my class ends in 2 hours and i need to pass my class! :'(

OpenStudy (squirrels):

Ok so it says to find the surface area of a "sphere" This is the S.A. of a sphere \[4\pi r^2\] \[\pi=3.14\] r=radius and the 2 \[?^2\] is an exponent. Which means you multiply that number by the number by "2".

OpenStudy (tkhunny):

I don't really follow your logic. If you are "seriously bad at math", why do you think you should pass? I do not believe you are seriously bad at math. Based on your urgency, maybe you just let it sit around too long? You can't learn it in five minutes. It sometimes take time, particularly when you are SURE you can't do it. Stop trying to get someone else to do your homework or take your exam. You either know it or you don't if there remains only two hours. Time to pay the piper, perhaps.

OpenStudy (squirrels):

that just an example leah, im just showing you what an exponent is. that little 2. for example, \[4^2=4\times4\]

OpenStudy (anonymous):

I know where you are coming from but i have faced hardships that caused me to get behind and since i was little i have not really been able to comprehend math

OpenStudy (squirrels):

Anyways the diameter is 76 cm and the radius=half of the diameter can you tell me the radius?

OpenStudy (anonymous):

38

OpenStudy (squirrels):

good now lets go back to the surface area formula of a sphere. \[4\pi r^2\] now we can plug in the radius. \[4\pi 38^2\] Do you know what to do now?

OpenStudy (anonymous):

i got 1444

OpenStudy (squirrels):

ok good now it looks like this \[4\pi1444\] Now what?

OpenStudy (anonymous):

im not too sure what goes next to be honest

OpenStudy (squirrels):

Ok thats fine. What we do next is multiply. \[\pi=3.14\] \[4\times3.14\times1444\]

OpenStudy (squirrels):

Im sure what you know what to do next :p

OpenStudy (anonymous):

181364.64

OpenStudy (anonymous):

is that correct?

OpenStudy (squirrels):

Yup thats correct! (: good job. medal for you

OpenStudy (anonymous):

but its not an option for the answers

OpenStudy (anonymous):

Find the surface area of a sphere with a diameter of 76 cm. 3850 π cm2 1925 π cm2 5776 π cm2 7701 π cm2 these are the options

OpenStudy (squirrels):

OHHH thats right, well we just take a few steps back. \[4\pi r^2\] \[4\pi 38^2\] \[4\pi 1444\] \[5776\pi\]

OpenStudy (squirrels):

its basically the same thing we just did, but this time we did multiply the equation with \[\pi\]

OpenStudy (squirrels):

didnt* not did

OpenStudy (anonymous):

so how do i know which is the answer

OpenStudy (squirrels):

i have a feeling youre still not getting it

OpenStudy (anonymous):

sorry im just nervous!

OpenStudy (anonymous):

i have to finish my class before 12 and its gonna be 12 soon

OpenStudy (squirrels):

well look at the explanation i just gave you :p it shows you what to do

OpenStudy (anonymous):

can i ask you another question?

OpenStudy (squirrels):

yes you may

OpenStudy (anonymous):

Quadrilateral RSTU has vertices R (-3, 1) , S (-2, 3) , T (2, 3) , and U (3, 1). If you translate the quadrilateral four units down, what are the vertices of R'S'T'U'? R' (-3, -3) , S' (-2, -1) , T '(2, -1) , and U' (3, -3) R' (1, 1) , S' (2, 3) , T '(6, 3) , and U '(7, 1) R' (-7, 1) , S' (-6, 3) , T '(-2, 3) , and U '(-1, 1) R (-3, 3) , S (-2, 1) , T (2, 1) , and U (3, 3)

OpenStudy (tkhunny):

"since i was little i have not really been able to comprehend math" I have heard this over and over and over. It is almost always a false belief. Maybe it is true in your case, but I have not seen it in your questions. At the moment, I don't believe you are stuck with it. However, you seem oddly immune to the idea of one question per thread. Post on a new thread, please.

OpenStudy (squirrels):

Yes, please post a new thread, Leah

OpenStudy (anonymous):

why do i have to post a new one?

OpenStudy (squirrels):

Because users may get help from this thread, but its not gonna help with differenet questions all over the same thread :p

OpenStudy (anonymous):

ill make a new one

OpenStudy (tkhunny):

Single Questions in a Single Thread are MUCH more likely to be useful to someone else. It is very difficult to index or search multiple discussions on a single thread.

OpenStudy (squirrels):

Danke

OpenStudy (squirrels):

Beat ya to it Tk ;)

OpenStudy (tkhunny):

I used bigger words. :-)

OpenStudy (squirrels):

haha touche